### v1

# Average Cost Optimality of Partially Observed MDPs: Contraction of Non-Linear Filters and Existence of Optimal Solutions

Abstract.
The average cost optimality is known to be a challenging problem for partially observable stochastic control, with few results available beyond the finite state, action, and measurement setup, for which somewhat restrictive conditions are available. In this paper, we present explicit and easily testable conditions for the existence of solutions to the average cost optimality equation where the state space is compact. In particular, we present a contraction based analysis, which is new to the literature to our knowledge, building on recent regularity results for non-linear filters.
Key words and phrases:
Non-linear filtering, unique ergodicity, filter stability
1. Introduction
We study optimal control for partially observable Markov decision processes (PODMPs), under the average cost criterion. Let $\mathbb{X}$ denote a standard Borel space which is the state space of a partially observed controlled Markov process. Let $\mathbb{B}(\mathbb{X})$ be its Borel $\sigma$-field. Let $\mathbb{C}_{b}(\mathbb{X})$ be the set of all continuous, bounded functions on $\mathbb{X}$. Here and throughout the paper $\mathbb{Z}_{+}$ denotes the set of non-negative integers and $\mathbb{N}$ denotes the set of positive integers. Let $\mathbb{Y}$ be a standard Borel space denoting the observation space of the model, and let the state be observed through an observation channel $Q$. The observation channel, $Q$, is defined as a stochastic kernel (regular conditional probability) from $\mathbb{X}$ to $\mathbb{Y}$, such that $Q(\,\cdot\,|x)$ is a probability measure on the power set $P(\mathbb{Y})$ of $\mathbb{Y}$ for every $x\in\mathbb{X}$, and $Q(A|\,\cdot\,):\mathbb{X}\to[0,1]$ is a Borel measurable function for every $A\in P(\mathbb{Y})$. A decision maker (DM) is located at the output of the channel $Q$, and hence it only sees the observations $\{Y_{t},\,t\in\mathbb{Z}_{\geq 0}\}$ and chooses its actions from $\mathbb{U}$, the action space which is a compact set. An admissible policy $\gamma$ is a sequence of control functions $\{\gamma_{t},\,t\in\mathbb{Z}_{\geq 0}\}$ such that $\gamma_{t}$ is measurable with respect to the $\sigma$-algebra generated by the information variables $I_{t}=\{Y_{[0,t]},U_{[0,t-1]}\},\quad t\in\mathbb{N},\quad I_{0}=\{Y_{0}\},$ where
(1) $U_{t}=\gamma_{t}(I_{t}),\quad t\in\mathbb{Z}_{\geq 0},$
[…]
$\Pr\bigl{(}(X_{0},Y_{0})\in B\bigr{)}=\int_{B}\mu(dx_{0})Q(dy_{0}|x_{0}),\quad B\in\mathcal{B}(\mathbb{X}\times\mathbb{Y}),$
where $\mu$ is the (prior) distribution of the initial state $X_{0}$, and
$\Pr\biggl{(}(X_{t},Y_{t})\in B\,\bigg{|}\,(X,Y,U)_{[0,t-1]}=(x,y,u)_{[0,t-1]}\biggr{)}$
$=\int_{B}\mathcal{T}(dx_{t}|x_{t-1},u_{t-1})Q(dy_{t}|x_{t}),$
$B\in\mathcal{B}(\mathbb{X}\times\mathbb{Y}),t\in\mathbb{N},$ where $\mathcal{T}$ is the transition kernel of the model which is a stochastic kernel from $\mathbb{X}\times\mathbb{U}$ to $\mathbb{X}$.
We let the agent’s goal be to minimize the expected discounted cost of its actions over an infinite period of time.
$J_{\beta}(\mu,\gamma)=E_{\mu}^{\gamma}\left[\sum_{t=0}^{\infty}\beta^{t}c\left(X_{t},U_{t}\right)\right]$
for some discount factor $\beta\in(0,1)$, over the set of admissible policies $\gamma\in\Gamma$, where $c:\mathbb{X}\times\mathbb{U}\rightarrow\mathbb{R}$ is the stage-wise measurable cost function, and the expectation $E_{\mu}^{\mathcal{T},\gamma}$ is taken over the initial state probability measure $\mu$ and the transition kernel $\mathcal{T}$ under policy $\gamma$. The optimal cost for the discounted infinite horizon is defined as
$J_{\beta}^{*}(\mu)=\inf_{\gamma\in\Gamma}J_{\beta}(\mu,\gamma).$
The average cost control problem under partial observations involves finding the optimal policy that minimizes the average cost of the system over an infinite horizon:
$J^{*}(\mu)=\inf_{\gamma\in\Gamma}J(\mu,\gamma)$
where
$J(\mu,\gamma)=\limsup_{n\to\infty}\frac{1}{n}E_{\mu}^{\gamma}\left[\sum_{t=0}^{\infty}c\left(X_{t},U_{t}\right)\right].$
1.1. Notation and preliminaries
It is known that any POMDP can be reduced to a completely observable Markov process [1], [2], whose states are the posterior state distributions or ’beliefs‘ of the observer; that is, the state at time $t$ is
$z_{t}(\,\cdot\,):=P\{X_{t}\in\,\cdot\,|y_{0},\ldots,y_{t},u_{0},\ldots,u_{t-1}\}\in{\mathcal{P}}(\mathbb{X}).$
We call this equivalent process the filter process . We denote by $\mathcal{Z}:={\mathcal{P}}(\mathbb{X})$ the set of probability measures on $(\mathbb{X},\mathbb{B}(\mathbb{X}))$ under the weak topology, where, under this topology $\mathcal{Z}$ is also a standard Borel space. The filter process has state space $\mathcal{Z}$ and action space $\mathbb{U}$. Let $\mathbb{P}(\mathcal{Z})$ denote the probability measures on $\mathcal{Z}$, equipped with the weak topology.
The transition probability $\eta$ of the filter process can be determined via the following equation [3, 2, 1]:
(2) $\eta(\cdot\mid z,u)=\int_{\mathbb{Y}}1_{\{F(z,u,y)\in\cdot\}}P(dy\mid z,u),$
where
$P(\cdot\mid z,u)=\operatorname{Pr}\left\{Y_{n+1}\in\cdot\mid Z_{n}=z,U_{n}=u\right\}$
from $\mathcal{Z}\times\mathbb{U}$ to $\mathbb{Y}$ and
$F(z,u,y):=\operatorname{Pr}\left\{X_{n+1}\in\cdot\mid Z_{n}=z,U_{n}=u,Y_{n+1}=y\right\}$
from $\mathcal{Z}\times\mathbb{U}\times\mathbb{Y}$ to $\mathcal{Z}$.
The one-stage cost function $\tilde{c}:\mathcal{Z}\times\mathbb{U}\rightarrow\mathbb{R}$ is a Borel measurable function and it is given by
$\tilde{c}(z,u):=\int_{\mathbb{X}}c(x,u)z(dx),$
where $c:\mathbb{X}\times\mathbb{U}\rightarrow\mathbb{R}$ is the stage-wise cost function.
This way, we obtain a completely observable Markov decision process from the POMDP, with the components $(\mathcal{Z},\mathbb{U},\tilde{c},\eta)$. For the filter process, the information variables is defined as
$\tilde{I}_{t}=\left\{Z_{[0,t]},U_{[0,t-1]}\right\},\quad t\in\mathbb{N},\quad\tilde{I}_{0}=\left\{Z_{0}\right\}$
Let $\{\mu_{n},\,n\in\mathbb{N}\}$ be a sequence in $\mathcal{P}(\mathbb{X})$. The sequence $\{\mu_{n}\}$ is said to converge to $\mu\in\mathcal{P}(\mathbb{X})$ weakly if
(3) $\int_{\mathbb{X}}f(x)\mu_{n}(dx)\to\int_{\mathbb{X}}f(x)\mu(dx)$
for every continuous and bounded $f:\mathbb{X}\to\mathbb{R}$.
For two probability measures $\mu,\nu\in\mathcal{P}(\mathbb{X})$, the total variation metric is given by
$\|\mu-\nu\|_{TV}:=$ $2\sup_{B\in{\mathcal{B}}(\mathbb{X})}|\mu(B)-\nu(B)|$
(4) $=$ $\sup_{f:\,\|f\|_{\infty}\leq 1}\bigg{|}\int f(x)\mu(dx)-\int f(x)\nu(dx)\bigg{|},$
where the supremum is over all measurable real $f$ such that $\|f\|_{\infty}=\sup_{x\in\mathbb{X}}|f(x)|\leq 1$.
Finally, the bounded-Lipschitz metric $\rho_{BL}$ [4, p.109] can also be used to metrize weak convergence:
(5) $\rho_{BL}(\mu,\nu)=\sup_{\|f\|_{BL}\leq 1}\biggl{|}\int_{\mathbb{X}}f(e)\mu(de)-\int_{\mathbb{X}}f(e)\nu(de)\biggr{|},$
where
$\|f\|_{BL}:=\|f\|_{\infty}+\sup_{e\neq e^{\prime}}\frac{f(e)-f(e^{\prime})}{d_{\mathbb{X}}(e,e^{\prime})},$
and $d_{\mathbb{X}}$ is the metric on $\mathbb{X}$.
When $\mathcal{X}$ is compact, one way to metrize $\mathcal{Z}$ under the weak convergence topology is via the Kantorovich-Rubinstein metric (also known as the Wasserstein metric of order $1$) ([5], Theorem 8.3.2) defined as follows
(6) $W_{1}(\mu,\nu):=\sup\left\{\int_{\mathbb{X}}f(x)\mu(dx)-\int_{\mathbb{X}}f(x)\nu(dx):f\in\operatorname{Lip}(\mathbb{X},1)\right\},$
$\mu,\nu\in\mathcal{Z}$, where for $k\in\mathbb{N}$,
$\operatorname{Lip}(\mathbb{X},k)=\{f:\mathbb{X}\to\mathbb{R},\;\left\lVert f\right\rVert_{L}\leq k\}.$
Definition 1.1.
[[6], Equation 1.16.] For a kernel operator $K:S_{1}\rightarrow\mathcal{P}\left(S_{2}\right)$ we define the Dobrushin coefficient as:
$\delta(K)=\inf\sum_{i=1}^{n}\min\left(K\left(x,A_{i}\right),K\left(y,A_{i}\right)\right)$
where the infimum is over all $x,y\in S_{1}$ and all partitions $\left\{A_{i}\right\}_{i=1}^{n}$ of $S_{2}$.
1.2. Average Cost Optimality Equation
Average cost optimality equation (ACOE) plays a crucial role for the analysis and the existence results of MDPs under the infinite horizon average cost optimality criteria. The triplet $(h,\rho^{*},\gamma^{*})$, where $h,\gamma:\mathcal{Z}\to\mathds{R}$ are measurable functions and $*\rho*\in\mathds{R}$ is a constant forms the ACOE if
$h(z)+\rho^{*}$ $=\inf_{u\in\mathds{U}}\left\{\tilde{c}(z,u)+\int h(z_{1})\eta(dz_{1}|z,u)\right\}$
(7) $=\tilde{c}(z,\gamma^{*}(z))+\int h(z_{1})\eta(dz_{1}|z,\gamma^{*}(z))$
for all $z\in\mathcal{Z}$. It is well known that (see e.g. [3, Theorem 5.2.4]) if (1.2) is satisfied with the triplet $(h,\rho^{*},\gamma^{*})$, and furthermore if $h$ satisfies
$\sup_{\gamma\in\Gamma}\lim_{t\to\infty}\frac{E_{z}^{\gamma}[h(Z_{t})]}{t}=0,\quad\forall z\in\mathcal{Z}$
then $\gamma^{*}$ is an optimal policy for the POMDP under the infinite horizon average cost optimality criteria, and
$J^{*}(z)=\inf_{\gamma\in\Gamma}J(z,\gamma)=\rho^{*}\quad\forall z\in\mathcal{Z}.$
We will refer to the function $h$, the relative value function, for the rest of the paper. Note that there may not be a unique relative value function $h$ that satisfies the ACOE, however, any $h$ that satisfies the ACOE can be used for optimality analysis.
1.3. Statement of the Main Result
Now we state the main result of our paper.
Assumption 1.
  1. (1)
    $(\mathbb{X},d)$ is a bounded compact metric space with diameter $D$ (where $D=\sup_{x,y\in\mathbb{X}}d(x,y)$).
  2. (2)
    The transition probability $\mathcal{T}(\cdot\mid x,u)$ is continuous in total variation in $(x,u)$, i.e., for any $\left(x_{n},u_{n}\right)\rightarrow(x,u),\mathcal{T}\left(\cdot\mid x_{n},u_{n}\right)\rightarrow\mathcal{T}(\cdot\mid x,u)$ in total variation.
  3. (3)
    There exists $\alpha\in R^{+}$such that
    $\left\|\mathcal{T}(\cdot\mid x,u)-\mathcal{T}\left(\cdot\mid x^{\prime},u\right)\right\|_{TV}\leq\alpha d(x,x^{\prime})$
    for every $x,x^{\prime}\in\mathbb{X}$, $u\in\mathbb{U}$.
  4. (4)
    There exists $K_{1}\in\mathbb{R}^{+}$ such that
    $|c(x,u)-c(x^{\prime},u)|\leq K_{1}d(x,x^{\prime}).$
    for every $x,x^{\prime}\in\mathbb{X}$, $u\in\mathbb{U}$.
  5. (5)
    The cost function $c$ is bounded and continuous.
  6. (6)
    $K_{2}:=\frac{\alpha D(3-2\delta(Q))}{2}<1.$
We thus state the following main theorem.
Theorem 1.1.
Under Assumption 1, a solution to the average cost optimality equation (ACOE) exists. This leads to the existence of an optimal control policy, and optimal cost is constant for every initial state.
As can be seen, the testability/verification of these criteria is not overly demanding. We provide two examples to illustrate this, one in a continuous setting and the other in a discrete setting.
1.4. Examples
The first example will be for the discrete case. For the case with finite $\mathbb{X}$, consider the discrete metric $d$ defined as follows:
$d\left(x,x^{\prime}\right)=\begin{cases}1&\text{ if }x\neq x^{\prime}\\ 0&\text{ if }x=x^{\prime}.\end{cases}$
With this choice of metric, the diameter $D$ is equal to $1$.
Example 1.1.
Let $\mathbb{X}=\{0,1,2,3\}$, $\mathbb{Y}=\{0,1\}$, $\mathbb{U}=\{0,1\}$, $\epsilon\in(0,1/2)$, and let $c$ be any function from $X\times U$ to $R^{+}$. Now, consider the transition and measurement matrices given by:
$\mathcal{T}_{0}=\left(\begin{array}[]{cccc}1/2&1/3&1/6&0\\ 0&1/2&1/6&1/3\\ 1/2&1/6&0&1/3\\ 1/3&1/3&1/3&0\end{array}\right)\quad Q=\left(\begin{array}[]{ccc}3/4-\epsilon&1/4+\epsilon\\ 3/4-\epsilon&1/4+\epsilon\\ 1/4+\epsilon&3/4-\epsilon\\ 1/4+\epsilon&3/4-\epsilon\end{array}\right),$
$\mathcal{T}_{1}=\left(\begin{array}[]{cccc}1/3&1/2&1/6&0\\ 0&1/3&1/2&1/6\\ 1/2&1/3&0&1/6\\ 1/3&1/3&1/3&0\end{array}\right)$
Please note that $\delta(Q)>1/2$, and we can choose $\alpha=1$. Therefore, according to Theorem 1.1, the Average Cost Optimal Equation (ACOE) has an optimal policy.
The second example will be for the continuous state case.
Example 1.2.
Let $\mathbb{X}=[0,2]$, $\mathbb{Y}=\{0,1\}$, $\mathbb{U}=[0,12]$, let the transition kernel $\mathcal{T}(.\mid x,u)=U(0,min(2,1+((x+u)/7)))$, where $U$ stands for uniform distribution, let measurement kernel $Q(x)=\lfloor x\rfloor$ and let cost function $c(x,u)=x+u$.
For this example, we observe that $\delta(Q)=0$. Furthermore, considering any control input $u$ in the set $\mathbb{U}$ and any states $x$ and $x^{\prime}$ within the space $\mathbb{X}$ such that $x^{\prime}<x$, we can derive the following bound:
$\left\|T(\cdot|x,u)-T(\cdot|x^{\prime},u)\right\|_{TV}$
$\leq\left\|U(0,1+(x^{\prime}+u)/7)-U(0,1+(x+u)/7)\right\|_{TV}$
$=2\left(1-\frac{1+\frac{x^{\prime}+u}{7}}{1+\frac{x+u}{7}}\right)$
$=2\frac{x-x^{\prime}}{x+u+7}$
$<\frac{2}{7}d(x,x^{\prime}).$
For this example, we have $D=supd(x,x^{\prime})=2$, and by choosing $\alpha=\frac{2}{7}$ and $K_{1}=1$, we can set $K_{2}$ to $6/7$. This choice ensures that we satisfy the conditions outlined in Assumption 1. Consequently, we can directly apply Theorem 1.1 to this specific example.
Example 1.3.
Consider $\mathbb{X}=[0,1]$, $\mathbb{U}=[-p,p]$ and the cost function $c(x,u)=x-u$. Define the transition kernel $\mathcal{T}(.|x,u)=\bar{N}(x+u,\sigma^{2})$. Here, $\bar{N}(\mu,\sigma^{2})$ denote truncated version of $N(\mu,\sigma^{2})$, where the support is restricted to $[0,1]\subset\mathbb{R}$. Its probability density function $f$ is given by
$f(x;\mu,\sigma)=\frac{1}{\sigma}\frac{\varphi\left(\frac{x-\mu}{\sigma}\right)}{\Phi\left(\frac{1-\mu}{\sigma}\right)-\Phi\left(\frac{0-\mu}{\sigma}\right)}$
Here, $\varphi(\cdot)$ is the probability density function of the standard normal distribution and Φ()normal-$\Phi(\cdot)$ is its cumulative distribution function.
For any $0\leq x<y\leq 1$: $\frac{\lVert\mathcal{T}(.|y,u)-\mathcal{T}(.|x,u)\rVert_{TV}}{y-x}\leq\frac{\sqrt{2}}{\sigma\sqrt{\pi}}.$ This shows that the transition kernel $\mathcal{T}$ satisfies Assumption 1-3 with $\alpha=\frac{\sqrt{2}}{\sigma\sqrt{\pi}}$. If we choose $\sigma>3/\sqrt{2\pi}$, Assumption 1-6 is satisfied for every observation channel. Under this condition, by Theorem 1.1, the ACOE has a solution.
1.5. Literature Review and Comparison
In the following we present a detailed literature review and also make a comparison with out results. We will in particular highlight that many of the results in the literature are not easy to verify, which also serves to demonstrate that the problem is a challenging one.
References [7, 8, 9, 10] study the average-cost control problem under the assumption that the state space is finite; they provide reachability type conditions for the belief kernels. Reference [10] also provides a detailed literature review. An additional line of argument, via simulation and coupling, was introduced in [11, 12, 13, 14] considers the finite model setup and [13] considers the case with finite-dimensional real-valued state spaces under several technical assumptions on the controlled state process via a coupling argument. Under certain strong continuity requirements of the transition kernel, [15] extends the positivity condition to Polish spaces. We also note that, via the weak Feller property of non-linear filters, the convex analytic methods can also be utilized [16, Theorem 1.2], though the dependence on the initial condition is a limitation.
Borkar and Budhiraja’s [13] assumptions and results can be explained as follows. They consider $\mathbb{X}$, $\mathbb{Y}$, and $\mathbb{U}$ as Polish spaces, where $\mathbb{X}$ is a finite-dimensional Euclidean space (not necessarily compact, unlike ours) and $\mathbb{U}$ is a compact space. In their setup, unlike ours, the observation measurement depends on the previous time’s state and action. For $(x,u)\in\mathbb{X}\times\mathbb{U}$, the transition probability measure $p(x,u,dz,dy)\in\mathscr{P}(\mathbb{X}\times\mathbb{Y})$ is defined. [13] assumes that $p$ is continuous in a strong sense: Let $\lambda$ denote the Lebesgue measure on $\mathbb{X}$ and let $\eta\in\mathscr{P}(\mathbb{Y})$, assume a density function $\varphi(x,u,z,y)$ on $\mathbb{X}\times\mathbb{U}\times\mathbb{X}\times\mathbb{Y}$ such that $p(x,u,dz,dy)=\varphi(x,u,z,y)\lambda(dz)\eta(dy)$, with $\varphi(\cdot)>0$. It is assumed that $\varphi(x,u,z,y)$ is continuous. Note that by Scheff’e’s lemma, this implies that the transition kernel as well as the measurement kernel are continuous in total variation. The update rule is then given as
$Pr\left(X_{n+1}\in A,Y_{n+1}\in A^{\prime}\mid X_{[0,n]},Y_{[0,n]},U_{[0,n]}\right)$
$=\int_{A^{\prime}}\int_{A}\varphi\left(X_{n},U_{n},z,y\right)\lambda(\mathrm{d}z)\eta(\mathrm{d}y).$
Furthermore, to facilitate a coupling argument, [13] additionally assumes
(8) $\int\bar{\varphi}\left(x,u,x^{\prime}\right)\left(\frac{\varphi\left(x,u,x^{\prime},y\right)}{\bar{\varphi}\left(x,u,x^{\prime}\right)}\right)^{1+\varepsilon_{0}}\lambda\left(dx^{\prime}\right)\eta(dy)<\infty$
for some $\varepsilon_{0}>0$, where $\bar{\varphi}(x,u,x^{\prime})=\int\varphi(x,u,x^{\prime},y)\eta(dy)$. Finally, the following Lyapunov-type assumption is assumed (which always holds when $\mathbb{X}$ is compact).
Assumption 2.
There exist $\mathscr{V}\in\mathbb{C}(\mathbb{X})$ and $\hat{\mathscr{V}}\in\mathbb{C}(\mathbb{X})$ satisfying $\lim_{\|x\|\rightarrow\infty}\mathscr{V}(x)=\infty$ and $\lim_{\|x\|\rightarrow\infty}\hat{\mathscr{V}}(x)=\infty$, and under any wide sense admissible $\left\{Z_{n}\right\}$,
$E\left[\hat{\mathscr{V}}\left(X_{n+1}\right)\mid\mathscr{F}_{n}\right]-\hat{\mathscr{V}}\left(X_{n}\right)\leqslant-\mathscr{V}\left(X_{n}\right)+\hat{C}I_{\hat{B}}\left(X_{n}\right),$
$\limsup_{n\rightarrow\infty}\frac{E\left[\hat{\mathscr{V}}\left(X_{n}\right)\right]}{n}=0,$
where $\hat{C}>0$ and $\hat{B}=\{x\in S:\|x\|\leqslant\hat{R}\}$ for some $\hat{R}>0$ and $\mathscr{F}_{n}=\sigma(X_{[0,n]},Y_{[0,n]},U_{[0,n]})$.
Theorem 1.2 (Theorem 4.1 of [13]).
For the system described above and under the mentioned assumptions as well as Assumption 2, a solution to the average cost optimality equation (ACOE) exists. This implies the existence of an optimal control policy, and the optimal cost remains constant for every initial state.
Compared to this result, we present complementary conditions which are explicit and testable. Additionally, for the compact setup, we do not have continuity assumptions the measurement kernels and the condition (8) is not needed. Our paper looks to be the first one where contraction properties of filter kernels is utilized.
In [7], Platzman addresses the finite-state average-cost control problem with finite state, observation, and action spaces, under very restrictive reachability, subrectangularity, and detectability conditions.
Runggaldier and Stettner in [9] consider $X$ and $Y$ as finite and $U$ as compact. They prove that under the following positivity condition [9], the Average Cost Optimal Equation (ACOE) has a bounded solution:
Assumption 3 ([9]).
$\inf_{i,j\in\boldsymbol{X}}\inf_{u,u^{\prime}\in\boldsymbol{U}}\inf_{\left\{C\in\mathbb{B}(\mathbb{X}):[T(C|i,u)]>0\right\}}\frac{T\left(C|j,u^{\prime}\right)}{T(C|i,u)}>0.$
We note that Example 1.1 does not meet this condition.
Borkar [11] employed a coupling argument for the finite case. In order for the Average Cost Optimal Equation (ACOE) to possess a solution, the following is assumed.
Assumption 4 ([11] Assumption A).
There exist constants $K_{0}\in\mathbb{R}^{+}$ and $\delta\in(0,1)$ such that $\sup_{i,j}\sup_{\gamma}P(\tau>n|X^{\prime}_{0}=i,X_{0}=j)\leq K_{0}\delta^{n}$ holds for all $n\geq 0$. This supremum is taken over all wide-sense admissible policies, and $\tau$ represents the coupling time, i.e., $\tau=\min\{n:X^{\prime}_{n}=X_{n}\}$.
However, verifying this assumption is not simple due to the necessity of taking the supremum over all wide-sense admissible policies (which is a strict superset of deterministic admissible policies).
Hsu, Chuang, and Arapostathis [10] consider a finite $\mathbb{X}$ and $\mathbb{Y}$ and a compact $\mathbb{U}$. They provide two different sets of assumptions under which the Average Cost Optimal Equation (ACOE) has a solution.
Assumption 5 (Assumption 2 of [10]).
$\mathcal{Z}_{\epsilon}=\{\mu\in\mathcal{Z}:\mu(x)>\epsilon\quad\forall x\in X\}$. There exist constants $\epsilon>0$, $k_{0}\in\mathbb{N}$, and $\alpha<1$ such that if $z_{*}(\beta)\in\arg\min_{z\in Z}J_{\beta}(z)$, then for each $\beta\in[\alpha,1)$ we have
$\max_{1\leq k\leq k_{0}}\mathbb{P}_{z_{*}(\beta)}^{\gamma_{\beta}}(Z_{k}\in\mathcal{Z}_{\epsilon})\geq\epsilon,$
where $Z_{k}$ is the filter process and $\gamma_{\beta}$ is the optimal policy for the $\beta$-discounted horizon cost problem.
Because checking this assumption can be quite complex, an alternative easier-to-verify assumption is provided
Assumption 6 ([10]).
There exist $k\geq 1$, and $\Delta>0$ such that, for all admissible $\{U_{t}\}$
$\mathbb{P}(X_{k}=j|X_{0}=i,U_{t-1},Y_{t},1\leq t\leq k)\geq\Delta\text{ for all }i,j\in X.$
For the second set of assumptions, the conditions are relaxed to some extent when the cost function $c$ is continuous.
Assumption 7 ([10]).
There exists $k\geqslant 1$, and $\Delta>0$ such that, for all $y^{k}\in\mathbb{Y}^{k}$ and $u^{k}\in\mathbb{U}^{k}$,
$P_{ij}\left(y^{k}|u^{k}\right)\geq\Delta\sum_{\ell\in\boldsymbol{X}}P_{\ell j}\left(y^{k}|u^{k}\right)\quad\forall i,j\in\boldsymbol{X}$
Theorem 1.3 (Theorem 11 of [10]).
Under either Assumption 5 or Assumption 7, the Average Cost Optimal Equation (ACOE) possesses a bounded solution.
Finally, Stettner, in [15], extends the results of [9] to Polish state spaces and observation spaces, encompassing both nondegenerate and degenerate observations: For $u_{1},u_{2}\in\mathbb{U}$ and $\mu,\nu\in P(\mathbb{X})$ define
$\lambda\left(u_{1},u_{2},\mu,\nu\right):=\inf_{\{C:T(C|\mu,u_{1})>0\}}\frac{T(C|\nu,u_{2})}{T(C|\mu,u_{1})}.$
Then define $\lambda(\mu,\nu)=\inf_{u\in U}\lambda(u,u,\mu,\nu)$.
Assumption 8 ([15]).
$\lambda\left(u^{1},u_{n}^{2},\mu,\nu_{n}\right)\rightarrow 1$ and $\lambda\left(u_{n}^{2},u^{1},\nu_{n},\mu\right)\rightarrow 1$ when $u_{n}^{2}\rightarrow u^{1}$ and $\nu_{n}\Rightarrow\mu$. Similarly, assume that both $\lambda\left(\nu_{n},\mu\right)$ and $\lambda\left(\mu,\nu_{n}\right)$ converge to 1 as $\nu_{n}\Rightarrow\mu$.
Theorem 1.4 (Theorem 5.6 of [15]).
Let $\mathbb{X}$ and $\mathbb{Y}$ be Polish spaces, $\mathbb{U}$ be a compact space, and $c$ be a continuous and bounded function. Under Assumptions 3 and 8, the Average Cost Optimality Equation (ACOE) admits a bounded solution.
With this review, we have both summarized some existing related work, and also highlighted that our paper presents accessible and testable conditions, compared with most of the literature reviewed above.
In our paper, we present a contraction based analysis for non-linear filters, which is a novel contribution and which we expect to have significant consequences for learning theoretic and approximation results.
2. Proof of the Main Theorem
Recently, [17] presented the following regularity results for controlled filter processes:
Theorem 2.1 ([17], Theorem 7-i).
Assume that $\mathbb{X}$ and $\mathbb{Y}$ are Polish spaces. If Assumption 1-3 is fulfilled, then we have
$\rho_{BL}\left(\eta(\cdot\mid z),\eta\left(\cdot\mid z^{\prime}\right)\right)\leq 3\left(1+\alpha\right)\rho_{BL}\left(z,z^{\prime}\right).$
Theorem 2.2 ([17], Theorem 7-iv).
$\rho_{BL}\left(\eta(\cdot|z,u),\eta(\cdot|z^{\prime},u)\right)\leq(3-2\delta(O))(1-\tilde{\delta}(\mathcal{T}))\|z-z^{\prime}\|_{TV}$
for any $z,z^{\prime}\in{\mathcal{P}}(\mathbb{X})$.
Building on these results and their proof method, the following result essentially follows from [18]. A proof sketch is given in the appendix.
Theorem 2.3.
[18] Assume that $\mathbb{X}$ and $\mathbb{Y}$ are Polish spaces. If Assumptions 1-1,3 are fulfilled, then we have
$W_{1}\left(\eta(\cdot\mid z_{0},u),\eta\left(\cdot\mid z_{0}^{\prime},u\right)\right)\leq\left(\frac{\alpha D(3-2\delta(Q))}{2}\right)W_{1}\left(z_{0},z_{0}^{\prime}\right).$
for all $z_{0},z_{0}^{\prime}\in\mathcal{Z}$, $u\in\mathbb{U}$.
As we assume that $\mathbb{X}$ is compact, under the $W_{1}$ (that is, the $1$-Wasserstein or Kantorovich-Rubinstein) metric, $\mathcal{Z}$ is compact.
Under Assumption 1-4, we have that $\tilde{c}$ is Lipschitz continuous, since
$|\tilde{c}(z,u)-\tilde{c}(z^{\prime},u)|$
(9) $=\left|\int_{\mathbb{X}}c(x,u)z(dx)-\int_{\mathbb{X}}c(x,u)z^{\prime}(dx)\right|\leq K_{1}W_{1}\left(z,z^{\prime}\right).$
Theorem 2.4.
[19, Theorem 2] Under Assumption 1-2, the transition probability $\eta(\cdot\mid z,u)$ of the filter process is weakly continuous in $(z,u)$.
We also highlight that additional recent results are available on the weak Feller property, such as those in [20]; however, as there are no restrictions on $Q$, this result [19, Thm.2] is relevant here.
Lemma 2.1.
Under Assumption 1, for any $\beta$ the value function $J_{\beta}^{*}$ is Lipschitz continuous with coefficient $K$, where $K=\frac{K_{1}}{1-\beta K_{2}}$.
Proof.
We adopt the approach in the proof of [21, Theorem 4.37]. Let $f\in\operatorname{Lip}(\mathcal{Z},k)$ for some $k>0$. Then $g=\frac{f}{k}\in\operatorname{Lip}(\mathcal{Z},1)$ and therefore for all $u\in\mathrm{U}$ and $z,y\in\mathcal{Z}$ we have
(10) $\left|\int_{\mathcal{Z}}f(x)\eta(dx\mid z,u)-\int_{\mathcal{Z}}f(x)\eta(dx\mid y,u)\right|$
(11) $=k\left|\int_{\mathcal{Z}}g(x)\eta(dx\mid z,u)-\int_{\mathcal{Z}}g(x)\eta(dx\mid y,u)\right|$
(12) $\leq kW_{1}(\eta(\cdot\mid z,u),\eta(\cdot\mid y,u))\leq kK_{2}W_{1}(z,y),$
by Theorem 2.3. Let $T$ be the Bellman optimality operator.
$(Tf)(x)=\min_{u\in U}\left\{c(x,u)+\beta\int_{y\in\mathcal{Z}}\eta(dy|x,u)f(y)\right\}$
We know that $J^{*}_{\beta}$ satisfy $TJ^{*}_{\beta}=J^{*}_{\beta}$ [3, Lemma 4.2.6]. $T$ is a contraction, so by Banach fixed point theorem $T^{n}f$ converges to $J^{*}_{\beta}$.
$\mid Tf(z)$ $-Tf(y)\mid$
$\leq\max_{u\in\mathrm{U}}\bigg{\{}|c(z,u)-c(y,u)|$
$+\beta\left|\int_{\mathbf{Z}}f(x)\eta(dx\mid z,u)-\int_{\mathbf{Z}}f(x)\eta(dx\mid y,u)\right|\bigg{\}}$
$\leq K_{1}W_{1}(z,y)+\beta kK_{2}W_{1}(z,y)=\left(K_{1}+\beta kK_{2}\right)W_{1}(z,y)=:M_{1}W_{1}(z,y).$
By induction we have for all $n\geq 2$
$T^{n}f\in\operatorname{Lip}\left(Z,M_{n}\right),$
where $M_{n}=K_{1}+\beta K_{2}M_{n-1}$ and thus $M_{n}=K_{1}\sum_{i=0}^{n-1}\left(\beta K_{2}\right)^{i}+k\left(\beta K_{2}\right)^{n}$. We have that $M_{n}\to\frac{K_{1}}{1-\beta K_{2}}$ since $K_{2}<1$. Hence, $T^{n}f\in\operatorname{Lip}\left(\mathcal{Z},\frac{K_{1}}{1-\beta K_{2}}\right)$ for all $n$, and therefore, $J_{\beta}^{*}\in\operatorname{Lip}\left(\mathcal{Z},\frac{K_{1}}{1-\beta K_{2}}\right)$ since it is closed with respect to the sup-norm. Taking $k\leq\frac{K_{1}}{1-\beta K_{2}}$, we certify that the fixed point satisfies the desired Lipschitz continuity.
For any $\beta\in(0,1)$, $J^{*}_{\beta}\in\operatorname{Lip}\left(\mathcal{Z},\frac{K_{1}}{1-K_{2}}\right)$. Therefore, for any $z_{0}\in\mathcal{Z}$
(13) $h_{\beta}(z)=J_{\beta}(z)-J_{\beta}(z_{0})\leq\frac{K_{1}}{1-K_{2}}W_{1}(z,z_{0})\leq\frac{K_{1}\cdot D}{1-K_{2}}.$
In view of the results above, we now introduce a crucial auxiliary theorem that will play a pivotal role in establishing our main result:
Assumption 9.
  1. (1)
    The one stage cost function c is bounded and continuous.
  2. (2)
    The stochastic kernel $\mathcal{T}(\cdot\mid x,u)$ is weakly continuous in $(x,u)\in\mathbb{X}\times\mathbb{U}$, i.e., if $\left(x_{k},u_{k}\right)\rightarrow(x,u)$, then $\mathcal{T}\left(\cdot\mid x_{k},u_{k}\right)\rightarrow$ $\mathcal{T}(\cdot\mid x,u)$ weakly.
  3. (3)
    $\mathbb{U}$ is compact.
  4. (4)
    $\mathbb{X}$ is $\sigma$-compact, that is, $\mathbb{X}=\cup_{n}S_{n}$ where $S_{n}\subset S_{n+1}$ and each $S_{n}$ is compact.
There exists $\alpha\in(0,1)$ and $N\geq 0$, and a state $z_{0}\in\mathbb{X}$ such that,
  1. (5)
    $-N\leq h_{\beta}(z)\leq N$ for all $z\in\mathbb{X}$ and $\beta\in[\alpha,1)$, where
    $h_{\beta}(z)=J_{\beta}(z)-J_{\beta}\left(z_{0}\right),$
    for some fixed $z_{0}\in\mathbb{X}$.
  2. (6)
    The sequence $\left\{h_{\beta(k)}\right\}$ is equicontinuous, where $\{\beta(k)\}$ is a sequence of discount factors converging to 1 which satisfies $\lim_{k\rightarrow\infty}(1-\beta(k))J_{\beta(k)}^{*}(z)=\rho^{*}$ for all $z\in\mathbb{X}$ for some $\rho^{*}\in[0,L]$.
Theorem 2.5.
[22, Theorem 7.3.3] Under Assumption 9, a solution to the average cost optimality equation (ACOE) exists, leading to the existence of an optimal control policy, and optimal cost is constant for every initial state.
A similar theorem, with slightly stronger continuity conditions (not applicable to our setup) under [3, Assumption 4.2.1], can be found in [3, Theorem 5.5.4]. Using this auxiliary result, we are now ready to present the proof of the main theorem.
Proof of Theorem 1.1.
In view of Lemma 2.1, all conditions of Theorem 2.5 are satisfied. Assumption 9-2 follows from Theorem 2.4, and Assumption 9-5 follows from (13). Notably, the equi-continuity condition, Assumption 9-6 holds as a result of the relation
$h_{\beta}(z)=J_{\beta}(z)-J_{\beta}(z_{0})\leq\frac{K_{1}}{1-K_{2}}W_{1}(z,z_{0})$
for all $\beta\in(0,1)$.
3. Concluding Remarks
The average cost optimality is a challenging problem, especially under the average cost criterion. In this paper, we presented explicit and easily testable conditions for the existence of solutions to the average cost optimality equation where the state space is compact. A comparison with the related literature and several examples are presented. Notably, our paper appears to be the first one where a contraction result is presented and utilized for the optimality analysis. Further research will involve providing implications of our contraction based analysis on approximations and learning theoretic results.
4. Appendix
4.1. Proof sketch of Theorem 2.3
Proof.
We equip $\mathcal{Z}$ with the metric $W_{1}$ to define the Lipschitz seminorm $\left\lVert f\right\rVert_{L}$ of any Borel measurable function $f:\mathcal{Z}\to\mathbb{R}$.
$W_{1}\left(\eta(\cdot\mid z_{0},u),\eta\left(\cdot\mid z_{0}^{\prime},u\right)\right)$
(14) $=\sup_{f\in\operatorname{Lip}(\mathcal{Z}),\;\left\lVert f\right\rVert_{\infty}\leq D/2}\bigg{|}\int_{\mathbb{Y}}f\left(z_{1}\left(z_{0}^{\prime},u,y_{1}\right)\right)P\left(dy_{1}\mid z_{0}^{\prime},u\right)$
(15) $-\int_{\mathbb{Y}}f\left(z_{1}\left(z_{0},u,y_{1}\right)\right)P\left(dy_{1}\mid z_{0},u\right)\bigg{|}$
For any $f:\mathcal{Z}\to\mathbb{R}$ such that $\left\lVert f\right\rVert_{L}\leq 1$ and $\left\lVert f\right\rVert_{\infty}\leq D/2$, we have
$\left|\int_{\mathbb{Y}}f\left(z_{1}\left(z_{0}^{\prime},u,y_{1}\right)\right)P\left(dy_{1}\mid z_{0}^{\prime},u\right)-\int_{\mathbb{Y}}f\left(z_{1}\left(z_{0},u,y_{1}\right)\right)P\left(dy_{1}\mid z_{0},u\right)\right|$
$\leq\left|\int_{\mathbb{Y}}f\left(z_{1}\left(z_{0}^{\prime},u,y_{1}\right)\right)P\left(dy_{1}\mid z_{0}^{\prime},u\right)-\int_{\mathbb{Y}}f\left(z_{1}\left(z_{0}^{\prime},u,y_{1}\right)\right)P\left(dy_{1}\mid z_{0},u\right)\right|$
$+\int_{\mathbb{Y}}\left|f\left(z_{1}\left(z_{0}^{\prime},u,y_{1}\right)\right)-f\left(z_{1}\left(z_{0},u,y_{1}\right)\right)\right|P\left(dy_{1}\mid z_{0},u\right)$
(16) $\leq\frac{D}{2}\left\|P\left(\cdot\mid z_{0}^{\prime},u\right)-P\left(\cdot\mid z_{0},u\right)\right\|_{TV}$
(17) $+\int_{\mathbb{Y}}\left|f\left(z_{1}\left(z_{0}^{\prime},u,y_{1}\right)\right)-f\left(z_{1}\left(z_{0},u,y_{1}\right)\right)\right|P\left(dy_{1}\mid z_{0},u\right).$
For the first term:
(18) $\left\|P\left(\cdot\mid z_{0}^{\prime},u\right)-P\left(\cdot\mid z_{0},u\right)\right\|_{TV}$
$=\sup_{\|g\|_{\infty}\leq 1}\left|\int g\left(y_{1}\right)P\left(dy_{1}\mid z_{0}^{\prime},u\right)-\int g\left(y_{1}\right)P\left(dy_{1}\mid z_{0},u\right)\right|$
(19) $\leq(1-\delta(Q))\left\|\mathcal{T}\left(dx_{1}\mid z_{0}^{\prime},u\right)-\mathcal{T}\left(dx_{1}\mid z_{0},u\right)\right\|_{TV}$
by Dobrushin [6].
(20) $\left\|\mathcal{T}\left(dx_{1}\mid z_{0}^{\prime},u\right)-\mathcal{T}\left(dx_{1}\mid z_{0},u\right)\right\|_{TV}$
$=\sup_{\|g\|_{\infty}\leq 1}\left(\int g\left(x_{1}\right)T\left(dx_{1}\mid z_{0}^{\prime},u\right)-\int g\left(x_{1}\right)T\left(dx_{1}\mid z_{0},u\right)\right)$
(21) $=\sup_{\|g\|_{\infty}\leq 1}\left(\int\tilde{g_{g}}(x_{0})z_{0}^{\prime}(dx_{0})-\tilde{g_{g}}(x_{0})z_{0}(dx_{0})\right)$
where
$\tilde{g_{g}}(x)=\int g\left(x_{1}\right)T\left(dx_{1}\mid x,u\right).$
For all $x_{0}^{\prime},x_{0}\in\mathbb{X}$, we have,
$\left\|\mathcal{T}\left(dx_{1}\mid x_{0}^{\prime},u\right)-\mathcal{T}\left(dx_{1}\mid x_{0},u\right)\right\|_{TV}\leq\alpha d(x_{0},x_{0}^{\prime}).$
As a result, we get $\tilde{g}/\alpha\in Lip(\mathbb{X})$.
Then, by inequality (18) and (20) we can write
(22) $\left\|P\left(\cdot\mid z_{0}^{\prime},u\right)-P\left(\cdot\mid z_{0},u\right)\right\|_{TV}\leq\alpha(1-\delta(Q))W_{1}\left(z_{0}^{\prime},z_{0}\right).$
Finally, we can analyze the second term in (17)
$\int_{\mathbb{Y}}\left|f\left(z_{1}\left(z_{0}^{\prime},u,y_{1}\right)\right)-f\left(z_{1}\left(z_{0},u,y_{1}\right)\right)\right|P\left(dy_{1}\mid z_{0},u\right)$
$\leq\int_{\mathbb{Y}}W_{1}(z_{1}\left(z_{0}^{\prime},u,y_{1}\right),z_{1}\left(z_{0},u,y_{1}\right))P\left(dy_{1}\mid z_{0},u\right)$
(23) $=\int_{\mathbb{Y}}\sup_{g\in\operatorname{Lip}(\mathbb{X})}\left(\int_{\mathbb{X}}g(x_{1})w_{y_{1}}(dx_{1})\right)P\left(dy_{1}\mid z_{0},u\right),$
where $w_{y_{1}}=\left(z_{1}\left(z_{0}^{\prime},u,y_{1}\right)-z_{1}\left(z_{0},u,y_{1}\right)\right)$ which is a signed measure on $\mathbb{X}$. By the Measurable Selection Theorem choose measurable $g_{y}\in\arg\sup_{g\in\operatorname{Lip}(\mathbb{X})}\left(\int_{\mathbb{X}}g(x)w_{y}(dx)\right).$
After that we can continue with the equation (4.1),
$\int_{\mathbb{Y}}\sup_{g\in\operatorname{Lip}(\mathbb{X})}\left(\int_{\mathbb{X}}g(x_{1})w_{y_{1}}(dx_{1})\right)P\left(dy_{1}\mid z_{0},u\right)$
(24) $=\int_{\mathbb{Y}}\int_{\mathbb{X}}g_{y_{1}}(x_{1})z_{1}\left(z_{0}^{\prime},u,y_{1}\right)(dx_{1})P(dy_{1}\mid z_{0},u)$
(25) $-\int_{\mathbb{Y}}\int_{\mathbb{X}}g_{y_{1}}(x_{1})z_{1}\left(z_{0}^{\prime},u,y_{1}\right)(dx_{1})P(dy_{1}\mid z_{0}^{\prime},u)$
(26) $+\int_{\mathbb{Y}}\int_{\mathbb{X}}g_{y_{1}}(x_{1})z_{1}\left(z_{0}^{\prime},u,y_{1}\right)(dx_{1})P(dy_{1}\mid z_{0}^{\prime},u)$
(27) $-\int_{\mathbb{Y}}\int_{\mathbb{X}}g_{y_{1}}(x_{1})z_{1}\left(z_{0},u,y_{1}\right)(dx_{1})P(dy_{1}\mid z_{0},u)$
For the first term, we can write by the same argument as earlier
$\left\lVert\int_{\mathbb{X}}g_{y_{1}}(x_{1})z_{1}\left(z_{0}^{\prime},u,y_{1}\right)(dx_{1})\right\rVert_{\infty}\leq\left\lVert g_{y_{1}}\right\rVert_{\infty}\leq D/2.$
So,
(28) $\int_{\mathbb{Y}}\int_{\mathbb{X}}g_{y_{1}}(x_{1})z_{1}\left(z_{0}^{\prime},u,y_{1}\right)(dx_{1})P(dy_{1}\mid z_{0},u)$
$-\int_{\mathbb{Y}}\int_{\mathbb{X}}g_{y_{1}}(x_{1})z_{1}\left(z_{0}^{\prime},u,y_{1}\right)(dx_{1})P(dy_{1}\mid z_{0}^{\prime},u)$
$\leq\frac{D}{2}\left\|P\left(\cdot\mid z_{0}^{\prime},u\right)-P\left(\cdot\mid z_{0},u\right)\right\|_{TV}$
(29) $\leq\alpha\frac{D}{2}(1-\delta(Q))W_{1}\left(z_{0}^{\prime},z_{0}\right)$
by inequality (22).
For the second term, we can write by smoothing
(30) $\int_{\mathbb{Y}}\int_{\mathbb{X}}g_{y_{1}}(x_{1})z_{1}\left(z_{0}^{\prime},u,y_{1}\right)(dx_{1})P(dy_{1}\mid z_{0}^{\prime},u)$
$-\int_{\mathbb{Y}}\int_{\mathbb{X}}g_{y_{1}}(x_{1})z_{1}\left(z_{0},u,y_{1}\right)(dx_{1})P(dy_{1}\mid z_{0},u)$
(31) $=\int_{\mathbb{X}}\omega(x_{1})\mathcal{T}\left(dx_{1}\mid z_{0}^{\prime},u\right)-\int_{\mathbb{X}}\omega(x_{1})\mathcal{T}\left(dx_{1}\mid z_{0},u\right)$
where
$\omega(x_{1})=\int_{\mathbb{Y}}g_{y_{1}}(x_{1})Q\left(dy_{1}\mid x_{1}\right).$
For any $x^{\prime},x^{\prime\prime}\in\mathbb{X}$,
$\int_{\mathbb{X}}\omega(x)\mathcal{T}\left(dx\mid x^{\prime\prime},u\right)-\int_{\mathbb{X}}\omega(x)\mathcal{T}\left(dx\mid x^{\prime},u\right)$
$\leq\left\lVert\omega\right\rVert_{\infty}\left\|\mathcal{T}(\cdot\mid x^{\prime\prime},u)-\mathcal{T}\left(\cdot\mid x^{\prime},u\right)\right\|_{TV}$
$\leq\left\lVert\omega\right\rVert_{\infty}\alpha d(x^{\prime\prime},x^{\prime})$
$\leq\alpha\frac{D}{2}d(x^{\prime\prime},x^{\prime})$
So, by definition of the $W_{1}$ norm (6), we have
$\int_{\mathbb{X}}\omega(x_{1})\mathcal{T}\left(dx_{1}\mid z_{0}^{\prime},u\right)$
$\qquad-\int_{\mathbb{X}}\omega(x_{1})\mathcal{T}\left(dx_{1}\mid z_{0},u\right)$
$=\int_{\mathbb{X}}\int_{\mathbb{X}}\omega(x_{1})\mathcal{T}\left(dx_{1}\mid x_{0},u\right)z^{\prime}_{0}(dx_{0})$
$\qquad-\int_{\mathbb{X}}\int_{\mathbb{X}}\omega(x_{1})\mathcal{T}\left(dx_{1}\mid x_{0},u\right)z_{0}(dx_{0})$
(32) $\leq\alpha\frac{D}{2}W_{1}\left(z_{0},z_{0}^{\prime}\right).$
So, by the inequalities (25), (27), (28), (30), (4.1) we get
$\int_{\mathbb{Y}}\left|f\left(z_{1}\left(z_{0}^{\prime},u,y_{1}\right)\right)-f\left(z_{1}\left(z_{0},u,y_{1}\right)\right)\right|P\left(dy_{1}\mid z_{0},u\right)$
(33) $\leq\alpha\frac{D}{2}(2-\delta(Q))W_{1}\left(z_{0}^{\prime},z_{0}\right).$
If we take the supremum of the equation over all $f\in Lip(\mathcal{Z})$, then by using the inequalities (17),(22), and (4.1), we can write:
(34) $W_{1}\left(\eta(\cdot\mid z_{0},u),\eta\left(\cdot\mid z_{0}^{\prime},u\right)\right)\leq\left(\frac{\alpha D(3-2\delta(Q))}{2}\right)W_{1}\left(z_{0},z_{0}^{\prime}\right)$